• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. Monaghan

    Log entry time 03:23:48 on January10,2005

    Entry number 138804

    This entry is a followup to: 138803

    keyword=left bogies reset

    After both the primary and alternative tech-on-call did not respond to paging, the run coordinator decided to come in and help us find the computer/crate to reset. We made an access into the hall, found the correct VME crate for the left bogies and reset it. We were then able to move the spectrometer in the hall successfully and the gui controls were working again in the counting house.